# An Efficient Global Optimization Algorithm with Adaptive Estimates of the Local Lipschitz Constants

### v3

@@ 47       $\mathcal{Q}_{3}\leftarrow\operatorname*{arg\,max}\{\mathbf{v}\}$
48       $q^{\star}_{3}=\operatorname*{arg\,min}_{q_{3}\in\mathcal{Q}_{3}}\{\mathbf{l}\}$
@@ 68             $\mathbf{x}\leftarrow$ local_optimizer($\mathbf{x}_{p_{k}}$)
74                   $\mathcal{C}_{k}\leftarrow\mathcal{C}_{k}\cup\{j\in\mathcal{I}_{k}:||\mathbf{x}_{p_{k}}-\mathbf{x}_{j_{k}}||\leq r\}$
@@ 86             $\mathbf{S}_{{i^{*}_{k}}+2}\leftarrow\mathbf{S}_{i^{\star}_{k}}$
(a) $k=0$
(b) $k=1$
(c) $k=2$
(d) $k=3$
5 Numerical Results
In this section, we conduct a comprehensive evaluation of the HALO algorithm compared to other GO algorithms. The evaluation is performed on a diverse set of test functions organized into three different benchmarks. Additionally, we conduct a parametric study on the parameter $\beta$ from Eq. 19 to provide general recommendations to users regarding its value. Finally, we show how to extract insightful information from the objective function using our approach.
5.1 Experimental Setup
5.1.1 Algorithms
The following algorithms are included in our numerical experiments:
  • $\textbf{HALO}_{\text{L-BFGS-B}}$ and $\textbf{HALO}_{\text{SDBOX}}$: These are two versions of HALO that are coupled with different local optimizers. $\text{HALO}_{\text{L-BFGS-B}}$ utilizes a quasi-Newton method for bound-constrained problems (L-BFGS-B) [45], while $\text{HALO}_{\text{SDBOX}}$ uses a derivative-free coordinate descent method with an Armijo-type linesearch for bound-constrained optimization [46] (SD-BOX). Both versions of HALO have a single hyperparameter, $\beta$, which is defined in Eq. 19. This parameter determines the minimum size a selected partition must have before a local search can be initiated from its centroid. For our experiments, we fix $\beta$ to $10^{-4}$. In $\text{HALO}_{\text{L-BFGS-B}}$, the gradients used in the Quasi-Newton code are computed using finite difference methods.
  • HLO: This is similar to $\text{HALO}_{\text{SDBOX}}$, but it does not use adaptive estimates of the local Lipschitz constants as defined in Eq. 5. Instead, the lower bounds are calculated based on the estimate of the global Lipschitz constant only. In other words, all local Lipschitz constants are fixed as $\tilde{L}_{i_{k}}=\max_{i_{k}\in\mathcal{I}k}||\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}||$. This allows us to compare the convergence acceleration achieved by estimating the local Lipschitz constants using only the global Lipschitz constant estimate.
  • DIRECT [21]: This is the classical DIRECT algorithm, with a fixed hyperparameter $\epsilon$ set to $10^{-4}|f_{\text{min}}|$, as suggested in [21].
  • DIRMIN [44]: an efficient hybridization of the DIRECT algorithm. It starts local optimizations from each potentially optimal hyperrectangle identified by DIRECT. We use the same derivative-free local optimizer as in $\text{HALO}_{\text{SDBOX}}$, which is the algorithm defined in [46]. DIRMIN has shown improved convergence speed compared to DIRECT on various test functions [44] and has shown to be among the best DIRECT-type approaches [53]. It also outperformed other derivative-free hybrid GO algorithms in a real-world hull form shape optimization problem [54].
  • CMA-ES [5]: a popular stochastic meta-heuristic approach known for its strong performance in various applications. It is considered one of the most competitive algorithms in its category [55]. We fix the hyperparameter $\sigma_{0}$ to be $1/4$ of the domain, as recommended by the author. Due to the stochastic nature of the algorithm, we run 20 independent experiments for each test function.
  • L-SHADE [47]: another popular stochastic method that improves upon the SHADE algorithm by incorporating linear population size reduction. L-SHADE has demonstrated strong performance in various benchmarks and optimization challenges. We run 20 independent experiments to account for the stochastic nature of the algorithm and ensure reliable results. We used the implementation freely available at  [^2] https://github.com/xKuZz/pyade.
@@ 5.1.2 Test Functions
    First Benchmark: it is constructed using a function generator described in [49]. We considered 100 test functions for each dimension $N$ ranging in $\{2,3,4,6,8,10\}$, resulting in a total of 600 different test functions. The function generator requires two hyperparameters: the number of stationary points and their smoothness. For this experiment, we uniformly sampled the number of stationary points $S$ from the range $1$ to $100$, and the smoothness parameter was chosen uniformly from the range $[2,3]$, following the approach in [56].
@@
    Second Benchmark: it consists of objective functions generated using the methodology described in [57]. The original implementation was in C programming language, but we rewrote it in Python for our experiments. Similarly to the first benchmark, we considered 100 test functions for each dimension $N$ ranging in $\{2,3,4,6,8,10\}$, resulting in a total of 600 different test functions. This benchmark involves several hyperparameters: the distance $A$ from the global minimizer to the vertex of the paraboloid, the size of the basin of attraction $B$ of the global minimizer, and the number of local minima $C$. To create a challenging benchmark, we randomly set $A$ in the range $[0.8,1)$, $B$ in the range $[0.1,0.2)$, and $C$ from 3 to 10. These parameter settings ensure that the global minimizer is far from the vertex of the paraboloid and has a small basin of attraction.
    (a) $A=0.83,B=0.12,\\ C=7$
    (b) $A=0.93,B=0.1,\\ C=8$
    (c) $A=0.86,B=0.1,\\ C=4$
@@
    Third Benchmark: it comprises popular test functions frequently used to evaluate the performance of both local and global optimization algorithms, such as the Rosenbrock, Beale, Michalewicz, and Hartmann functions, as defined in [58]. In total, we considered 191 different functions for this benchmark. When a test function is applicable to multiple dimensions $N$, we used the same function for different values of $N$ ranging in $\{2,3,4,6,8,10\}$. Therefore, the total number of experiments using this benchmark is 496. In cases where a function has its global minimum at the center of the domain, we randomly shifted the global minimizer within the domain.
@@
5.1.3 Stopping Criteria
@@
The second termination criterion is based on the maximum number of function evaluations. In our experiments, we fixed this limit to 30000, irrespective of the problem’s dimensionality. If an algorithm exceeds 30000 function evaluations without satisfying the criterion defined in Eq. 34, the run is considered to have failed.
These termination criteria provide a balance between achieving accurate solutions and limiting the computational effort. By combining the relative error criterion with a maximum number of function evaluations, we ensure that the algorithms converge within a reasonable computational budget.
@@ 5.2 Comparison Results and Discussion
In the following three subsections, we present and analyze the numerical results obtained by each algorithm for the benchmarks considered in this study.
Given the large number of objective functions in each benchmark, we will provide a summary of the results using operational characteristics [59], also known as performance profiles [60]. More in particular the operational characteristic is defined as follows
$c(\gamma)=\frac{\text{Number of problems solved in less than $\gamma$ functionevaluations}}{\text{Total number of problems}}$ (35)
[…]
@@ 5.2.1 First Benchmark Results
Table 1 presents a summary of the numerical results obtained for the first benchmark. In terms of both AUOC and the number of function evaluations, the best-performing algorithm is $\text{HALO}_{\text{L-BFGS-B}}$, while the CMA-ES algorithm appears to be the least effective.
Notably, there is no significant difference in performance between the two versions of HALO equipped with different local optimizers. However, an interesting observation is a difference in performance between $\text{HALO}_{\text{SDBOX}}$ and HLO. $\text{HALO}_{\text{SDBOX}}$ solves $14.9\%$ more problems than HLO and demonstrates a much faster convergence, as evident from the operational characteristic plot shown in Fig. 7(a). Additionally, HLO performs worse than DIRECT. Upon analyzing this discrepancy, we found that HLO often selects only two hyperrectangles based on the estimate of the global Lipschitz constant. These hyperrectangles have the minimum objective function value (selected apriori) and the hyperrectangle with the largest diagonal and lowest objective function value. Consequently, if the global minimum is not within these two partitions, the overall convergence of the HLO algorithm is compromised.
@@
Table 1: Resume of the numerical results for the first benchmark.
Numerical results
Algorithm AUOC Percentage solved Average function evaluations
$\text{HALO}_{\text{L-BFGS-B}}$ $0.731$ $80.6$ $2920.0$
$\text{HALO}_{\text{SDBOX}}$ $0.726$ $80.4$ $3010.5$
HLO $0.590$ $65.5$ $3050.1$
DIRMIN $0.707$ $81.3$ $3984.2$
DIRECT $0.636$ $73.1$ $4032.6$
CMA-ES $0.207$ $22.6$ ($1.4$) $2572.3$
L-SHADE $0.276$ $37.5$ $(1.6)$ $11139.6$
@@ 5.2.2 Second Benchmark Results
Table 2 summarizes the numerical results obtained in the second benchmark. In this case, the best performing algorithm is $\text{HALO}_{\text{SD-BOX}}$, as it achieves the highest AUOC and solves the highest percentage of problems. However, its performance is quite similar to $\text{HALO}_{\text{L-BFGS-B}}$. In contrast, the CMA-ES algorithm remains the least effective in this benchmark.
@@
It is also interesting to analyze the difference between using the local Lipschitz constant, as in $\text{HALO}_{\text{SDBOX}}$, or the global Lipschitz constant, like in HLO. Additionally, this benchmark confirms that using an estimate of the local Lipschitz constant for each partition leads to better performance. $\text{HALO}_{\text{SDBOX}}$ solves more problems than HLO and uses fewer function evaluations. When comparing the performance of DIRECT and its hybrid counterpart DIRMIN, it is evident that DIRMIN is significantly less effective than DIRECT. This is primarily because too many local searches start and end up at the vertex of the paraboloid.
It appears that CMA-ES encounters significant difficulty in solving any of the objective functions. Our results indicate that almost $100\%$ of the time, CMA-ES gets trapped at the vertex of the paraboloid. Also L-SHADE experiences some difficulty with this benchmark, as is only able to solve $11.4$ of the problems.
Table 2: Resume of the numerical results for the second benchmark.
Numerical results
Algorithm AUOC Percentage solved Average functions evaluations
$\text{HALO}_{\text{L-BFGS-B}}$ $0.403$ $45.4$ $3487.9$
$\text{HALO}_{\text{SDBOX}}$ $0.404$ $45.8$ $3615.4$
HLO $0.382$ $44.8$ $4446.0$
DIRMIN $0.196$ $24.8$ $6327.2$
DIRECT $0.325$ $38.6$ $4831.4$
CMA-ES $0.0$ $0.0$ --
L-SHADE $0.100$ $11.4$ $(0.8)$ $26979.8$
@@ 5.2.3 Third Benchmark Results
Table 3 summarizes the numerical results obtained in the third and final benchmarks. The $\text{HALO}_{\text{L-BFGS-B}}$ algorithm emerges as the best performer, both in terms of AUOC and number of function evaluations. Conversely, the DIRECT algorithm appears to be relatively less effective.
There are some noticeable performance differences between the two versions of HALO employing different local optimizers. The variant equipped with the gradient-based optimizer, $\text{HALO}_{\text{L-BFGS-B}}$, demonstrates greater efficiency with fewer function evaluations compared to $\text{HALO}_{\text{SDBOX}}$. While $\text{HALO}_{\text{SDBOX}}$ and HLO exhibit similar problem-solving capabilities, $\text{HALO}_{\text{SDBOX}}$ achieves the solutions with an average of $\approx 400$ fewer function evaluations than HLO. This confirms that also in this benchmark, utilizing an estimate of the local Lipschitz constant accelerates the algorithm’s convergence towards the global minimizers more effectively than relying solely on the global Lipschitz constant. This observation is supported by the operational characteristics shown in Fig. 7(c).
Table 3: Resume of the numerical results for the third benchmark.
Numerical results
Algorithm AUOC Percentage solved Average function evaluations
$\text{HALO}_{\text{L-BFGS-B}}$ $0.700$ $71.9$ $872.77$
$\text{HALO}_{\text{SDBOX}}$ $0.692$ $73.1$ $1676.25$
HLO $0.688$ $73.7$ $2098.3$
DIRMIN $0.679$ $73.7$ $2543.6$
DIRECT $0.605$ $65.3$ $2282.5$
CMA-ES $0.632$ $69.1$ $(1.1)$ $2660.4$
L-SHADE $0.540$ $72.3$ $(1.0)$ $13810.6$
Although DIRMIN significantly accelerates the convergence of DIRECT, it does not outperform HALO and HLO in terms of AUOC and function evaluations. This is likely because DIRECT-type approaches, based on potentially optimal hyperrectangles, tend to sample many of them, leading to extensive exploration of sub-optimal regions in the domain $\mathcal{D}$. CMA-ES performs better than the DIRECT algorithm but remains less competitive compared to the other approaches considered. The behavior of L-SHADE is particularly notable when observing the operational characteristic in Fig. 7(c). The algorithm requires a large number of function evaluations to solve a substantial percentage of problems, which explains the low AUOC value it obtained.
@@
(a) First benchmark.
(b) Second Benchmark.
(c) Third benchmark.
Figure 7: Operational characteristics for the three benchmarks.
(a) First benchmark.
(b) Second benchmark.
(c) Third benchmark.
Figure 8: Percentage of the problem solved varying the dimensionality $N$ for the three benchmarks.
(a) First benchmark.
(b) Second benchmark.
(c) Third benchmark.
Figure 9: Average number of functions evaluations varying the dimensionality $N$ for the three benchmarks.
(a) Hard functions.
(b) Simple functions.

### v4

@@ 47  $\mathcal{Q}_{3}\leftarrow\operatorname*{arg\,max}\{\mathbf{v}\}$
    48  $q^{\star}_{3}=\operatorname*{arg\,min}_{q_{3}\in\mathcal{Q}_{3}}\{\boldsymbol{\ell}\}$
@@ 68     $\mathbf{x}\leftarrow$ local_optimizer($\mathbf{x}_{p_{k}}$)
[…]
@@ 86     $\mathbf{S}_{{i^{*}_{k}}+2}\leftarrow\mathbf{S}_{i^{\star}_{k}}$
  • DIRMIN [liuzzi2016exploiting]: an efficient hybridization of the DIRECT algorithm. It starts local optimizations from each potentially optimal hyperrectangle identified by DIRECT. We use the same derivative-free local optimizer as in $\text{HALO}_{\text{SDBOX}}$, which is the algorithm defined in [lucidi2002-COA]. DIRMIN has shown improved convergence speed compared to DIRECT on various test functions [liuzzi2016exploiting] and has shown to be among the best DIRECT-type approaches [stripinis2022directgo, stripinis2023derivative, stripinis2024benchmarking]. It also outperformed other derivative-free hybrid GO algorithms in a real-world hull form shape optimization problem [dagostino2018-AIAA-MAO].
  • CMA-ES [hansen2001completely]: a popular stochastic meta-heuristic approach known for its strong performance in various applications. It is considered one of the most competitive algorithms in its category [hansen2010comparing]. We fix the hyperparameter $\sigma_{0}$ to be $1/4$ of the domain, as recommended by the author. Due to the stochastic nature of the algorithm, we run 20 independent experiments for each test function.
  • L-SHADE [tanabe2014improving]: another popular stochastic method that improves upon the SHADE algorithm by incorporating linear population size reduction. L-SHADE has demonstrated strong performance in various benchmarks and optimization challenges. We run 20 independent experiments to account for the stochastic nature of the algorithm and ensure reliable results. We used the implementation freely available at  [^2] https://github.com/xKuZz/pyade.
@@ 5.1.2 Test Functions
    First Benchmark: it is constructed using a function generator described in [schoen1993wide]. We considered 100 test functions for each dimension $N$ ranging in $\{2,3,4,6,8,10\}$, resulting in a total of 600 different test functions. The function generator requires two hyperparameters: the number of stationary points and their smoothness. For this experiment, we uniformly sampled the number of stationary points $S$ from the range $1$ to $100$, and the smoothness parameter was chosen uniformly from the range $[2,3]$, following the approach in [cassioli2013global].
@@
    Second Benchmark: it consists of objective functions generated using the methodology described in [gaviano2003algorithm]. The original implementation was in C programming language, but we rewrote it in Python for our experiments. Similarly to the first benchmark, we considered 100 test functions for each dimension $N$ ranging in $\{2,3,4,6,8,10\}$, resulting in a total of 600 different test functions. This benchmark involves several hyperparameters: the distance $A$ from the global minimizer to the vertex of the paraboloid, the size of the basin of attraction $B$ of the global minimizer, and the number of local minima $C$. To create a challenging benchmark, we randomly set $A$ in the range $[0.8,1)$, $B$ in the range $[0.1,0.2)$, and $C$ from 3 to 10. These parameter settings ensure that the global minimizer is far from the vertex of the paraboloid and has a small basin of attraction.
    (a) A=0.83,B=0.12,C=7A=0.83,B=0.12,\\ C=7
    (b) A=0.93,B=0.1,C=8A=0.93,B=0.1,\\ C=8
    (c) A=0.86,B=0.1,C=4A=0.86,B=0.1,\\ C=4
@@
    Third Benchmark: it comprises popular test functions frequently used to evaluate the performance of both local and global optimization algorithms, such as the Rosenbrock, Beale, Michalewicz, and Hartmann functions, as defined in [jamil2013survey]. In total, we considered 191 different functions for this benchmark. When a test function is applicable to multiple dimensions $N$, we used the same function for different values of $N$ ranging in $\{2,3,4,6,8,10\}$. Therefore, the total number of experiments using this benchmark is 496. In cases where a function has its global minimum at the center of the domain, we randomly shifted the global minimizer within the domain.
@@
5.1.3 Termination Criteria and Evaluation Metric
@@
The second termination criterion is based on the maximum number of function evaluations. In our experiments, we fixed this limit to 50000, irrespective of the problem’s dimensionality. If an algorithm exceeds 50000 function evaluations without satisfying the criterion defined in Eq. 34, the run is considered to have failed.
These termination criteria provide a balance between achieving accurate solutions and limiting the computational effort. In particular, the budget of 50,000 evaluations is generally considered sufficient in the context of real-world black-box problems, as using significantly more would be unrealistic.
To evaluate the performance of the algorithms given the presence of large number of objective functions in each benchmark, we will provide a summary of the results using operational characteristics [strongin2013global], also known as performance profiles [dolan2002benchmarking]. More in particular the operational characteristic is defined as follows
[…]
$\text{AUOC}=\frac{1}{\gamma_{\text{max}}}\int_{0}^{\gamma_{\text{max}}}c(\gamma)\,d\gamma$ (36)
@@ 5.2 Comparison Results and Discussion
The numerical results highlight the significant advantage of using estimates of the local Lipschitz constant for accelerating convergence to the global optimum. Algorithms like $\text{HALO}_{\text{L-BFGS-B}}$ and $\text{HALO}_{\text{SDBOX}}$, which incorporate local Lipschitz estimates, outperformed direct-based approaches and those relying on sampling potentially optimal hyperrectangles. These methods consistently achieved higher total AUOC values, demonstrating their ability to more efficiently solve complex optimization problems.
In comparison, popular metaheuristics such as CMA-ES and L-SHADE showed weaker performance, failing to match the effectiveness of the Lipschitz-based algorithms. Despite their established use in global optimization, these methods did not achieve the same level of consistency or efficiency across the benchmarks.
Overall, the results confirm that incorporating local Lipschitz constant estimates can significantly improve optimization performance, offering a clear advantage over traditional DIRECT-type and metaheuristic approaches.
Figure 7: Operational characteristics for all the test functions considered in this study.
Table 1: Summary of the numerical results for different optimization algorithms across three benchmark tests. The table presents the total AUOC for each algorithm, along with their performance in the first, second, and third benchmarks.
Numerical results
Algorithm Total AUOC AUOC
First benchmark Second benchmark Third benchmark
$\text{HALO}_{\text{L-BFGS-B}}$ $0.631$ $0.773$ $0.425$ $0.709$
$\text{HALO}_{\text{SDBOX}}$ $0.630$ $0.768$ $0.427$ $0.710$
HLO $0.573$ $0.620$ $0.412$ $0.710$
HPLOR $0.515$ $0.520$ $0.356$ $0.701$
DIRMIN $0.552$ $0.759$ $0.218$ $0.707$
DIRECT $0.552$ $0.684$ $0.356$ $0.629$
CMA-ES $0.270$ $0.220$ $0.0$ $0.655$
L-SHADE $0.324$ $0.317$ $0.103$ $0.600$
@@ 5.2.1 First Benchmark Results
According to Table 1 the best-performing algorithm is $\text{HALO}_{\text{L-BFGS-B}}$, while the CMA-ES algorithm appears to be the least effective.
Notably, there is no significant difference in performance between the two versions of HALO equipped with different local optimizers. However, an interesting observation is a difference in performance between $\text{HALO}_{\text{SDBOX}}$ and HLO. $\text{HALO}_{\text{SDBOX}}$ solves more problems than HLO and demonstrates a much faster convergence, as evident from the operational characteristic plot shown in Fig. 9(a). Additionally, HLO performs worse than DIRECT. Upon analyzing this discrepancy, we found that HLO often selects only two hyperrectangles based on the estimate of the global Lipschitz constant. These hyperrectangles have the minimum objective function value (selected apriori) and the hyperrectangle with the largest diagonal and lowest objective function value. Consequently, if the global minimum is not within these two partitions, the overall convergence of the HLO algorithm is compromised. The same is valid for the HPLOR algorithm.
@@ 5.2.2 Second Benchmark Results
In this benchmark, the best performing algorithm is $\text{HALO}_{\text{SD-BOX}}$, as it achieves the highest AUOC and solves the highest percentage of problems (see Table 1. However, its performance is quite similar to $\text{HALO}_{\text{SD-BOX}}$. In contrast, the CMA-ES algorithm remains the least effective in this benchmark.
@@
It is also interesting to analyze the difference between using the local Lipschitz constant, as in $\text{HALO}_{\text{SDBOX}}$, or the global Lipschitz constant, like in HLO. Additionally, this benchmark confirms that using an estimate of the local Lipschitz constant for each partition leads to better performance. $\text{HALO}_{\text{SDBOX}}$ solves more problems than HLO and uses fewer function evaluations.
We can also observe that HPLOR, which does not use any estimate of the Lipschitz constants, performs worse compared to HALO and HLO. This suggests that leveraging information about the Lipschitz constants, both local (HALO) and global (HLO), provides valuable insights that help accelerate convergence toward the global optimum for this set of functions
When comparing the performance of DIRECT and its hybrid counterpart DIRMIN, it is evident that DIRMIN is significantly less effective than DIRECT. This is primarily because too many local searches start and end up at the vertex of the paraboloid.
It appears that CMA-ES encounters significant difficulty in solving any of the objective functions. Our results indicate that almost $100\%$ of the time, CMA-ES gets trapped at the vertex of the paraboloid. Also L-SHADE experiences some difficulty with this benchmark, as is only able to solve $11.4$ % of the problems.
It should be noted that the GKLS generator was reimplemented in this work, and since the full parameter sets used in prior studies were not released, the generated test functions cannot be replicated exactly, making direct comparison with previous results infeasible.
@@ 5.2.3 Third Benchmark Results
The numerical results for the third benchmark indicate that $\text{HALO}_{\text{SD-BOX}}$ and HLO are the best-performing algorithm, according to Table 1. Conversely, the L-SHADE algorithm appears to be relatively less effective.
While $\text{HALO}_{\text{SDBOX}}$ and HLO exhibit similar problem-solving capabilities, $\text{HALO}_{\text{SDBOX}}$ achieves the solutions with fewer function evaluations than HLO (see Fig. 11(c). This confirms that, in this benchmark as well, utilizing an estimate of the local Lipschitz constant accelerates the algorithm’s convergence towards the global minimizers more effectively than relying solely on the global Lipschitz constant. This observation is supported by the operational characteristics shown in Fig. 9(c). Although DIRMIN significantly accelerates the convergence of DIRECT, it does not outperform HALO and HLO in terms of AUOC and function evaluations. This is likely because DIRECT-type approaches, based on potentially optimal hyperrectangles, tend to sample many of them, leading to extensive exploration of sub-optimal regions in the domain $\mathcal{D}$. CMA-ES performs better than the DIRECT algorithm but remains less competitive compared to the other approaches considered. The behavior of L-SHADE is particularly notable when observing the operational characteristic in Fig. 9(c). The algorithm requires a large number of function evaluations to solve a substantial percentage of problems, which explains the low AUOC value it obtained.
@@
5.3 Assessing Variable Importance and Interpretability with HALO
As discussed in Section 4.5.1, HALO offers more than just an optimization routine; it also provides valuable insights into the importance of each variable in the objective function. After HALO completes its search, we can compute the variable importance using the matrix $\mathbf{G}$, which contains the approximate gradients $\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}$. By averaging these gradients as shown in Equation 33, we identify the directions in which the objective function is most sensitive.
In Fig. 8, we provide an illustrative example of the variable importance for three different objective functions: Eggholder, Dixon Price, and Adjiman respectively in Fig. 8(a), 8(b), and 8(c). By examining the variable importance defined in Eq. 33, we can observe that it accurately identifies the directions in which the objective function exhibits the most significant variation.
(a) Egg Holder.
(b) Dixon Price.
(c) Adjiman.